what expression raised to the fourth power is 81x^12y^8z^16
Answer by checkley71(8403)

(3 x~3 y~2 z~4)~4=81 x~12 y~8 z~16
3~4=3*3*3*3=81
X~3*X~3*X~3*X~3=X~12
Y~2*Y~2*Y~2*Y~2=Y~8
Z~4*Z~4*Z~4*Z~4=Z~16
THUS WE HAVE 81 X~12 Y~8 Z~16
HOPE THIS HELPS.